Determine the molar solubility of FeF2 in a solution that contains a buffer set to a pH of 2.35.
Ksp = 2.4 x 10-6
Ka for HF = 6.8 x 10-4
Answer #1

Solubility product of FeF2 = 2.4 × 10^ (-6)

The reaction of FeF2 is,

FeF2 -> Fe (2+) + 2 F(-)

Then Ksp = [ Fe 2+ ] × [ F - ] ^(-2)

2.4 * 10 ^ (-6) = X * X^(2)

2.4 * 10 ^ (-6) = X ^(3)

Or the value of X is cube root of 2.4 * 10^(-6) = 1.3338 which is the molar solubility.

So the value of solubility is 1.3338.